OpenStudy (anonymous):

find the derivative of: y=e^x/x^2-x

OpenStudy (anonymous):

quotient rule

OpenStudy (anonymous):

what is the alue of g(x), i know it cant be 0.

OpenStudy (anonymous):

value not alue

OpenStudy (anonymous):

by using uv' - vu'/v^2.... e^x( 2x-1) - (x^2-x.e^x)/ (x^2-x)^2 =e^x2x-e^x - e^x.x^2+e^x.x/ (x^2-x)^2

OpenStudy (anonymous):

omg, im so dumb, im sorry didnt realize g(x) = x^2-x

OpenStudy (anonymous):

im sorry

OpenStudy (anonymous):

thank you so much for your help!

OpenStudy (anonymous):

let f(x)= \[e^x\] let g(x)= \[x^2-x\] now, use this formula \[{f \prime x [g(x)]- [f(x)]g \prime x}/[g(x)]^2\]

OpenStudy (anonymous):

yes sorry, i just forgot that g=x^2-x .. thank you tho

OpenStudy (anonymous):

np